OpenStudy (anonymous):

How do i solve this?: 2 sin^2y = 5 -2 cos^2y

OpenStudy (kc_kennylau):

Use the identity \(\large\sin^2\theta+\cos^2\theta=1\) :) @douglaswinslowcooper will lead you :)

OpenStudy (anonymous):

Recall that sin^2 + cos^2 = 1

OpenStudy (anonymous):

Unfortunately, that leads to 2 = 5, wrong. Re-check question.

OpenStudy (kc_kennylau):

@missannemarie10 no solution BTW welcome to OS :D